A child sleds down a frictionless hill with vertical drop h. At the bottom is a level stretch where the coefficient of friction is 0.27.

If she slides 19 m across the level stretch, what's h? Express your answer with the appropriate units.
Answer #1

kinetic energy of child at bottom = potentail energy at top

= m*g*h

Workdone by friction = chnage in kinetic energy on frictional surface

Friction*d*cos(180) = KEF - KEi


-mue_k*N*d = 0 - m*g*h

-mue_k*m*g*d = -m*g*h

==> h = mue_k*d

= 0.27*19

= 5.13 m <<<<<<<-------------Answer